# CamReasoner: Reinforcing Camera Movement Understanding via Structured Spatial Reasoning

### v1

@@ 3.2.2 Reinforcement Learning Framework
The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight (set to $0.5$). The overall reward is formulated as:
@@ 4.1 Setup
We evaluate our model on the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks. To ensure visual grounding, each question is paired with two videos yielding opposite answers. During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128. This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.
@@
The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision. We employ a rollout batch size of 128 and a global batch size of 32. To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution. The total reward is defined as $R=r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$. Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
@@ 4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs in average accuracy.
@@
LLaVA-OneVision-7B [li2024llavaonevision] 46.8 13.5 12.6 16.9 23.7 20.2 10.7 14.4 33.5 33.6 16.9 31.4 19.3 20.8 18.8 22.2
LLaVA-Video-7B [zhang2025llavavideo] 54.7 15.2 16.5 19.3 27.1 23.6 16.2 16.9 33.6 36.8 26.9 37.2 16.1 21.7 22.1 25.6
InternVideo2-Chat-8B [wang2024internvideo2] 69.9 18.5 19.3 17.6 17.9 23.4 12.2 10.4 22.6 22.7 17.2 22.8 19.6 16.4 20.2 22.0
Tarsier-Recap-7B [yuan2025tarsier2] 59.7 15.1 25.7 23.7 28.8 21.5 14.4 15.0 22.8 27.3 24.6 21.6 15.2 18.7 30.7 21.0
InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 49.0 10.6 11.4 10.4 14.6 10.6 11.8 16.5 14.3 13.9 14.7 17.5 11.7 18.1 21.8 16.5
InternVL2.5-8B [chen2025expandingperformance] 67.9 12.9 28.1 25.9 23.4 23.2 18.6 32.1 37.4 30.9 37.6 36.9 11.5 25.3 23.4 29.5
InternVL2.5-26B [chen2025expandingperformance] 63.6 11.8 21.1 23.6 27.2 19.4 21.8 31.6 42.5 38.3 44.9 43.6 14.3 18.2 25.1 29.8
GPT-4o [openai2024gpt4ocard] 66.3 29.2 21.1 38.2 38.0 21.9 41.7 39.3 44.7 42.1 43.6 35.5 24.0 28.7 32.0 36.4
InternVL3-8B [zhu2025internvl3e] 61.2 15.5 18.8 29.0 30.5 27.3 29.5 28.1 41.6 49.3 42.0 36.5 21.3 22.3 20.1 31.5
InternVL3-78B [zhu2025internvl3e] 72.0 18.2 19.6 32.5 33.8 29.4 26.4 33.4 47.2 53.5 47.8 40.3 27.6 25.0 22.6 36.8
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 56.0 14.9 18.7 30.5 34.5 27.6 29.8 43.4 62.7 66.7 54.5 34.1 18.8 24.2 19.8 35.8
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 57.6 16.4 20.2 32.1 36.0 29.2 31.4 45.0 64.3 68.2 56.0 35.6 20.2 25.7 21.2 37.3
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 58.0 16.8 20.6 32.5 36.4 29.5 31.7 45.4 64.7 68.6 56.4 36.0 20.6 26.1 21.6 37.7
Previous SOTA
Cam-Motion-7B [lin2025towards] 83.9 38.6 27.8 47.8 67.9 50.0 54.5 75.8 79.2 83.8 76.3 67.6 32.3 41.0 73.6 60.0
Cam-Motion-32B [lin2025towards] 85.6 40.1 29.3 49.4 69.6 51.5 56.0 77.3 80.7 85.4 77.9 69.2 33.9 42.7 75.4 61.6
Cam-Motion-72B [lin2025towards] 86.8 41.3 30.5 50.6 70.7 52.6 57.1 78.5 81.9 86.6 79.1 70.4 35.0 43.8 76.6 62.8
@@
As presented in Tab. 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models. Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o (36.4%) and Qwen2.5-VL-72B (37.7%) but also yields a substantial improvement of 15.6% over the task-specific SFT baseline. This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks. Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds. Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the accuracy of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines. By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting. Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
Table 2: VQA evaluation on the CameraBench. We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions. Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
@@
mPLUG-Owl3-7B 51.8 64.9 61.5 48.6 49.2 54.1 53.2 45.9 63.4 55.8
LLaVA-Video-7B [zhang2025llavavideo] 53.5 66.1 57.2 52.1 49.9 54.9 59.9 51.3 68.0 58.8
LLaVA-OneVision-7B [li2024llavaonevision] 54.3 63.8 69.0 53.1 55.4 60.9 60.7 43.3 52.3 57.1
InternVideo2-Chat-8B [wang2024internvideo2] 52.4 64.4 51.7 50.2 49.7 52.2 48.5 50.9 50.6 51.3
Tarsier-Recap-7B [yuan2025tarsier2] 51.8 62.8 50.5 49.8 49.0 51.5 47.8 50.2 49.8 50.6
InternVL2.5-8B [chen2025expandingperformance] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
InternVL2.5-26B [chen2025expandingperformance] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
InternVL3-8B [zhu2025internvl3e] 54.4 59.8 57.5 51.3 49.7 58.1 55.2 50.0 50.0 54.5
InternVL3-78B [zhu2025internvl3e] 56.2 63.5 60.8 53.8 51.2 60.3 58.4 52.5 53.6 57.2
GPT-4o [openai2024gpt4ocard] 55.8 52.6 61.2 58.1 53.3 64.1 51.7 42.1 61.9 59.0
Gemini-2-Flash [comanici2025gemini25] 53.6 46.8 56.6 44.5 41.1 46.5 46.5 39.2 63.8 51.8
Gemini-2.5-Pro [comanici2025gemini25] 58.2 51.3 60.1 48.9 45.7 52.3 49.7 42.8 64.5 54.7
@@ 4.5 Ablation Study
Table 4: Ablation study of the reward weight $\lambda$. Values represent the Accuracy (%) across different evaluation tasks. A balanced weight of $\lambda=0.1$ achieves the optimal performance.

### v2

@@ 3.2.2 Reinforcement Learning Framework
The final reward $R(o)$ is a weighted interpolation between accuracy and format adherence. Let $\lambda$ denote the format weight. The overall reward is formulated as:
@@ 4.1 Setup
We evaluate our model using data from the CameraBench [lin2025towards] dataset, focusing on binary classification of motion primitives and the 10K-sample VQA subset covering 9 skills and 81 sub-tasks. We tested the accuracy of multiple existing MLLMs using the questions and videos across the datasets. To ensure visual grounding, each question is paired with two videos yielding opposite answers. During inference, we standardize video input to a maximum of 64 frames and 262,144 pixels ($128\times 32\times 32$) sampled at 1 FPS, with a batch size of 128. This setup evaluates the model across object-centric motion, scene dynamics, and complex logical verification.
@@
The RL phase is implemented using the verl framework with a learning rate of $2\times 10^{-6}$ and bf16 mixed-precision. We employ a rollout batch size of 128 and a global batch size of 32. To ensure stability, we incorporate an online filtering mechanism that prunes samples in the top and bottom 1% of the reward distribution. The total reward is defined as $R=(1-\lambda)\cdot r_{acc}+\lambda\cdot r_{fmt}$, with $\lambda=0.1$. Crucially, the RL policy is initialized from the first SFT epoch to provide a foundational reasoning grasp while maintaining sufficient entropy for exploration.
@@ 4.2 Quantitative Results
Table 1: Binary classification performance on camera movement understanding. CamReasoner-7B achieves state-of-the-art results across most categories, significantly outperforming SfM/SLAM baselines and general-purpose MLLMs.
@@
LLaVA-OneVision-7B [li2024llavaonevision] 48.2 13.9 13.0 17.4 24.4 20.8 11.0 14.8 34.5 34.6 17.4 32.3 19.9 21.4 19.4 22.9
LLaVA-Video-7B [zhang2025llavavideo] 56.3 15.7 17.0 19.9 27.9 24.3 16.7 17.4 34.6 37.9 27.7 38.3 16.6 22.4 22.8 26.4
InternVideo2-Chat-8B [wang2024internvideo2] 72.0 19.1 19.9 18.1 18.4 24.1 12.6 10.7 23.3 23.4 17.7 23.5 20.2 16.9 20.8 22.7
Tarsier-Recap-7B [yuan2025tarsier2] 61.5 15.6 26.5 24.4 29.7 22.1 14.8 15.5 23.5 28.1 25.3 22.2 15.7 19.3 31.6 21.6
InternLMXComposer2.5-7B [zhang2024internlmxcomposer25] 50.5 10.9 11.7 10.7 15.0 10.9 12.2 17.0 14.7 14.3 15.1 18.0 12.1 18.6 22.5 17.0
InternVL2.5-8B [chen2025expandingperformance] 69.9 13.3 28.9 26.7 24.1 23.9 19.2 33.1 38.5 31.8 38.7 38.0 11.8 26.1 24.1 30.4
InternVL2.5-26B [chen2025expandingperformance] 65.5 12.2 21.7 24.3 28.0 20.0 22.5 32.5 43.8 39.4 46.2 44.9 14.7 18.7 25.9 30.7
GPT-4o [openai2024gpt4ocard] 68.3 30.1 21.7 39.3 39.1 22.6 43.0 40.5 46.0 43.4 44.9 36.6 24.7 29.6 33.0 37.5
InternVL3-8B [zhu2025internvl3e] 63.0 16.0 19.4 29.9 31.4 28.1 30.4 28.9 42.8 50.8 43.3 37.6 21.9 23.0 20.7 32.4
InternVL3-78B [zhu2025internvl3e] 74.2 18.7 20.2 33.5 34.8 30.3 27.2 34.4 48.6 55.1 49.2 41.5 28.4 25.8 23.3 37.9
Qwen2.5-VL-7B [bai2025qwen25vltechnicalreport] 57.7 15.3 19.3 31.4 35.5 28.4 30.7 44.7 64.6 68.7 56.1 35.1 19.4 24.9 20.4 36.9
Qwen2.5-VL-32B [bai2025qwen25vltechnicalreport] 59.3 16.9 20.8 33.1 37.1 30.1 32.3 46.4 66.2 70.2 57.7 36.7 20.8 26.5 21.8 38.4
Qwen2.5-VL-72B [bai2025qwen25vltechnicalreport] 59.7 17.3 21.2 33.5 37.5 30.4 32.7 46.8 66.6 70.7 58.1 37.1 21.2 26.9 22.2 38.8
SOTA Baseline
Cam-Motion-7B [lin2025towards] 86.4 40.5 29.2 50.2 71.3 52.5 57.2 79.6 83.2 87.1 80.1 71.0 33.9 43.1 77.3 63.0
Cam-Motion-32B [lin2025towards] 88.2 42.1 30.8 51.9 73.1 54.1 58.8 81.2 84.7 88.8 81.8 72.7 35.6 44.8 79.2 64.7
Cam-Motion-72B [lin2025towards] 89.4 43.4 32.0 53.1 74.2 55.2 60.0 82.4 86.0 90.1 83.1 73.9 36.8 46.0 80.4 65.9
@@
As presented in Tab. 4.2, our proposed CamReasoner-7B demonstrates a significant performance leap over a broad spectrum of baseline models, ranging from traditional geometric frameworks to state-of-the-art large vision-language models. Collectively, our model achieves a superior overall accuracy of 78.4%, which not only dramatically outperforms general-purpose benchmarks like GPT-4o and Qwen2.5-VL-72B but also yields a substantial improvement over the task-specific SFT baseline. This performance gap underscores the model’s exceptional mastery of camera-centric reasoning, particularly in complex rotation and translation tasks. Specifically, CamReasoner-7B establishes new SOTA results in challenging categories such as Move Up (84.9%), Tilt Up (87.6%), and Pan Left (86.1%), suggesting that it has developed a robust spatial-temporal awareness capable of decoupling ego-motion from dynamic scene backgrounds. Furthermore, the model exhibits remarkable resilience in traditionally difficult scenarios like Move Out and Zooming, where perspective transformations are most aggressive; for instance, it elevates the performance of the Move Out category to 57.7%, nearly doubling the performance of several competitive baselines. By comparing these results with SFT-only variants, it becomes evident that while supervised learning provides the fundamental paradigm for motion identification, our reinforcement learning phase effectively sharpens the policy’s decision boundaries and maintains high policy entropy during the initial stages to avoid overfitting. Consequently, these results validate that our framework successfully bridges the gap between raw visual perception and high-level camera movement reasoning, providing a robust foundation for automated cinematography analysis.
Table 2: VQA evaluation on the CameraBench. We compare CamReasoner-7B against various state-of-the-art MLLMs across nine fine-grained dimensions with questions on camera movement type in given video clips. Our model consistently achieves the highest accuracy in all categories, demonstrating superior reasoning capabilities in camera movement understanding.
@@
mPLUG-Owl3-7B 53.1 66.4 62.9 49.8 50.4 55.4 54.5 47.1 64.9 57.2
LLaVA-Video-7B [zhang2025llavavideo] 54.8 67.7 58.6 53.4 51.1 56.3 61.3 52.6 69.6 60.2
LLaVA-OneVision-7B [li2024llavaonevision] 55.7 65.3 70.6 54.4 56.8 62.4 62.2 44.4 53.6 58.5
InternVideo2-Chat-8B [wang2024internvideo2] 53.7 65.9 53.0 51.5 51.0 53.5 49.7 52.2 51.8 52.6
Tarsier-Recap-7B [yuan2025tarsier2] 53.1 64.3 51.7 51.0 50.2 52.8 49.0 51.4 51.0 51.8
InternVL2.5-8B [chen2025expandingperformance] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL2.5-26B [chen2025expandingperformance] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
InternVL3-8B [zhu2025internvl3e] 55.8 61.2 58.9 52.6 50.9 59.5 56.5 51.2 51.2 55.8
InternVL3-78B [zhu2025internvl3e] 57.6 65.0 62.3 55.1 52.4 61.8 59.8 53.8 54.9 58.6
GPT-4o [openai2024gpt4ocard] 57.2 53.9 62.7 59.5 54.6 65.6 53.0 43.1 63.4 60.4
Gemini-2-Flash [comanici2025gemini25] 54.9 47.9 58.0 45.6 42.1 47.6 47.6 40.2 65.3 53.1
Gemini-2.5-Pro [comanici2025gemini25] 59.6 52.5 61.5 50.1 46.8 53.6 50.9 43.9 66.1 56.1
@@ 4.5 Ablation Study
Table 4: Ablation study of the reward weight $\lambda$. A balanced weight of $\lambda=0.1$ achieves the optimal performance.